OpenStudy (anonymous):

When trying to evaluate the indefinite integral of (2x * e^x), why does using the substitution rule give me an incorrect answer? (2x * e^x) u = e^x du/dx = e^x dx = du/e^x 2 * integral of [ x*e^x du/e^x ] e^x cancels out 2 * integral of [x] answer is 2 Again ... why is this wrong?

OpenStudy (anonymous):

because this is not a substitution method. you'll have to use integration by parts...

OpenStudy (anonymous):

I know that much. I am wondering why the substitution method does not work in this case.

OpenStudy (anonymous):

because as you can see, when you set u=e^x, there is still the factor of e^x when you calculate du... but the integrand only has an x (2x in this case)

OpenStudy (anonymous):

Why can't I just cancel it out? 2 * integral of [x * e^x du/e^x] It looks like it works (obviously it doesn't).

OpenStudy (anonymous):

last step answer is not 2. 2 * integral of [x] is x^2

OpenStudy (anonymous):

Put u=e^x as you mentioned, then du=e^x*dx . And again everything should be in the function of U. So,take u=e^x Taking log on both sides,we get log(u)=log(e^x) log(u)=xlog(e) log(u)=x (loge=10 Coming back to the equation integral of [2log(u)*u] Using integral by parts we get the answer as xe^x-(e^x)/2

OpenStudy (anonymous):

yes the do cancel but you'll then be left with integral x du. now since you're integrating wrt u, you'll need to change that x to lnu. so you have integral lnu du. even if you could do this new integral remember you need to switch back to the variabe x.